OpenStudy (chrisplusian):

Determine the base of the numbers in each case for the following operations to be correct: a) 14/2=5, 54/4=13, 24+17=40.

OpenStudy (chrisplusian):

sorry that should have been: a)14/2=5 b)54/4=13 c)24+17=40 I have the answers I need to know how to reason this out